OpenStudy (bananas1234):

9u^2 - 12u 3u Divide.

OpenStudy (anonymous):

\[\frac{ 9u^2-12u }{ 3u }? \implies \frac{ 9u^2 }{ 3u }-\frac{ 12u }{ 3u }\]

OpenStudy (gorv):

\[\frac{ 9u^2-12u }{ 3u }\]

OpenStudy (gorv):

take u out of numerator

OpenStudy (gorv):

\[\frac{ u*(9u-12) }{ 3*u }\]

OpenStudy (gorv):

take 3 out of numerator

OpenStudy (gorv):

\[\frac{ 3*u(3u-4) }{ 3u }\]